OpenStudy (anonymous):

Use a common denominator to determine whether the ratios form a proportion. 6 : 15 and 30 : 80 A. Yes. They form a proportion. B. No. They do not form a proportion.

OpenStudy (anonymous):

I think its B.

OpenStudy (anonymous):

cause 6/15 = 2/5 30/80 = 3/8

OpenStudy (anonymous):

obviously not the same so they are not proportional

OpenStudy (anonymous):

Am i right?

OpenStudy (anonymous):

@mahmood786123

OpenStudy (anonymous):

Im going with B.